# Contrastive Retrieval Heads Improve Attention-Based Re-Ranking

### v1

@@ 1 Introduction
To better characterize the functional role of attention heads, Wu et al. (2025) identify a subset of retrieval heads in transformers that specialize in extracting relevant information from long contexts. Their approach tracks copy–paste operations in Needle-in-a-Haystack tasks, which constrains its applicability to narrow question-answering scenarios. Building on this, Zhang et al. (2025) propose QR heads (query-focused retrieval heads), selected according to the absolute attention allocated to the correct answer. While aggregated QR-head scores improve performance on downstream retrieval tasks, the selection criterion overlooks relative ranking. For instance, a head may assign high attention to the gold document yet allocate even greater weight to irrelevant documents, undermining its discriminative capacity. Since the QR-head method does not penalize such cases, it may fail to identify the heads that are most effective at distinguishing relevant from irrelevant information.
@@
  • We demonstrate that CoRe heads identified using a small subset of the Natural Questions (NQ) dataset generalize across the BEIR benchmark (Thakur et al., 2021) and even across languages in the MLDR benchmark (Chen et al., 2024). Attention-based re-rankers using CoRe heads consistently outperform strong baselines, achieving state-of-the-art list-wise re-ranker performance.
  • We observe that top-scoring CoRe heads are primarily concentrated in the middle transformers layers. Exploiting this insight, we show that pruning most final layers incurs negligible loss in re-ranking accuracy, while reducing memory usage by $40\%$ and inference latency by $20\%$. This highlights the efficiency of CoRe heads for real-world retrieval systems.
@@ 2 Related Work
Zero-Shot Re-Ranking. Zero-shot re-ranking methods with LLMs can be broadly grouped into three main approaches: point-wise, pair-wise, and list-wise. Point-wise methods (Sachan et al., 2022; Liang et al., 2022) score each document independently, typically through logits or direct generation. These approaches are computationally efficient but often underperform due to the absence of cross-document comparison. Pair-wise methods (Qin et al., 2024) compare two documents at a time and aggregate over all pairs to form a ranking. While improving accuracy, the computational cost grows quadratically with the number of documents. List-wise methods (Ma et al., 2023; Sun et al., 2023; Chen et al., 2025) instead consider all documents jointly to produce a single ranked list. Although list-wise methods require long-context modeling, advances in LLMs with extended context windows (Chen et al., 2023; Jin et al., 2024a; Fu et al., 2024) have made this increasingly practical. Consequently, list-wise approaches now achieve both scalability and strong effectiveness. Our work follows this trajectory, focusing on attention-based list-wise re-ranking (Chen et al., 2025), which offers more efficiency and stability compared to generation-based methods (Ma et al., 2023; Sun et al., 2023).
Role of Attention Heads. A growing body of work in mechanistic interpretability examines how the activation of attention heads shapes model behavior. Michel et al. (2019) and Voita et al. (2019) show that only a small fraction of heads are necessary for translation tasks. Olsson et al. (2022) identify induction heads that capture repeated input patterns, and later extended the work to heads supporting in-context learning (Yin & Steinhardt, 2025; Ren et al., 2024). Other studies reveal specialized heads responsible for knowledge conflict (Shi et al., 2024; Jin et al., 2024b) and context distraction (Zhu et al., 2025). More recently, Wu et al. (2025) identify retrieval heads, which copy answer tokens from long contexts into the output, while Zhang et al. (2025) propose query-focused retrieval heads (QR heads) that better capture query–context interactions beyond direct copying. Collectively, these studies suggest that different subsets of heads specialize in distinct computational roles, and isolating the ones relevant to retrieval remains an open challenge. In this work, we build on the notion of retrieval heads (Wu et al., 2025; Zhang et al., 2025) as the mechanisms responsible for retrieving relevant information from long-context inputs – directly aligning with the re-ranking objective.
@@ 3 Background
Given a query $q$ and a set of $k$ candidate documents $D=\{d_{1},d_{2},\dots,d_{k}\}$ returned by a base retriever, the goal of the re-ranker is to reorder $D$ so that the documents are sorted in the order of relevance to $q$ . To achieve this, Chen et al. (2025) propose In-Context Re-ranking (ICR), an attention-based approach that leverages the attention signal from all layers and heads of a LLM to compute fine-grained relevance scores.
Given a LLM with $L$ layers and $H$ heads, ICR constructs an input prompt consisting of the list of $k$ documents followed by the query $q$. For each document $d_{i}$, ICR computes the relevance score $s_{i}$ by aggregating the attention score that each token in $d_{i}$ gets from $q$ across all heads and layers. The relevance score $s_{d_{i,j};q}$ of the $j$-th token in document $d_{i}$ with respect to the query $q$ is:
$s_{d_{i,j};q}=\frac{1}{|\mathcal{T}_{q}|}\sum_{l=1}^{L}\sum_{h=1}^{H}\sum_{t\in\mathcal{T}_{q}}a_{j,t}^{l,h},$ (1)
where $\mathcal{T}_{q}$ denotes the set of the query tokens, and $a_{j,t}^{l,h}$ denotes the attention score from the $t$-th token in query $q$ to the $j$-th token in document $d_{i}$ by the $h$-th attention head in the $l$-th layer.
To reduce bias relative to document length and low-information tokens (e.g., punctuation), ICR applies contextual calibration on the relevance score with a content-free query $q_{cf}$ (e.g. ‘N/A’) following Zhao et al. (2021). Specifically, the calibrated relevance score $s_{d_{i,j}}$ of the $j$-th token in document $d_{i}$ is then calculated as:
$s_{d_{i,j}}=s_{d_{i,j};q}-s_{d_{i,j};q_{cf}}.$ (2)
After eliminating tokens with abnormally negative calibrated scores, the relevance score of document $d_{i}$ is computed as the sum of remaining token scores.
@@
As described above, ICR computes relevance scores by aggregating attention signals from all heads and layers. While comprehensive, this aggregation can introduce redundancy and noise, lowering re-ranking performance. Recent work on QR heads (Zhang et al., 2025) demonstrates that attention from as few as $1\%$ of heads can improve retrieval effectiveness, indicating that many heads are unnecessary. A head $h$ is considered QR head if its attention to the gold document, quantified as $s_{gold}^{h}$, is the highest among all heads. Let $\mathcal{T}_{gold}$ be the set of tokens in the gold document, then $s_{gold}^{h}$ is computed as:
$s_{gold}^{h}=\frac{1}{|\mathcal{T}_{q}|}\sum_{t\in\mathcal{T}_{q}}\sum_{j\in\mathcal{T}_{gold}}a_{j,t}^{h}.$ (3)
However, the QR scoring criterion considers only absolute attention to relevant documents, without accounting for their relative ranking within each head’s distribution. This omission can result in the selection of misleading heads. As a consequence, QR-based re-ranking shows inconsistently performance across datasets: as shown in Figure 2, selecting the top $8$ QR heads reduces accuracy on the Quora duplicate-question retrieval task with Mistral 7B and Llama-3.1 8B, underperforming the ICR baseline that aggregates signals from all heads.
@@ 4 Contrastive Retrieval Heads
Addressing the limitations of prior approaches, we propose a contrastive head detection framework that identifies a small subset of Contrastive Retrieval heads (CoRe heads) specialized for document re-ranking. For clarity, we denote the default attention-based re-ranker as ICR, the QR-head re-ranker as QR-R, and the CoRe-head re-ranker as CoRe-R. Central to our method is a new scoring metric, $S_{CoRe}$, which measures a head’s ability to highlight the gold document while simultaneously suppressing attention to irrelevant ones. By selecting high-quality heads using this contrastive metric, CoRe-R achieves more consistent and robust improvements in re-ranking, outperforming ICR and QR-R on Quora (Figure 2) and across the BEIR benchmark, as demonstrated in Section 5.2.
@@
Our retrieval scoring metric follows the InfoNCE loss function (Oord et al., 2018), but requires no additional training. We consider the input prompt consisting of instructions (ignored in attention score computation), $k$ documents and a query $q$. Suppose the attention score from the $t$-th token in query $q$ to the $j$-th token in document $d$ by the an attention head $h$ is $a_{j,t}^{h}$, then the aggregated attention score head $h$ pays toward document $d$ is computed as:
$s_{d}^{h}=\frac{1}{|\mathcal{T}_{q}|}\sum_{t\in\mathcal{T}_{q}}\sum_{j\in\mathcal{T}_{d}}a_{j,t}^{h}.$ (4)
Let $s_{pos}^{h}$ be the attention score head $h$ pays toward the positive document, and $s_{neg,i}^{h}$ be the attention score head $h$ pays toward the $i$-th negative document. Then, the retrieval score of head $h$ is:
$S_{CoRe}(h)=\frac{\exp(s_{pos}^{h}/t)}{\exp(s_{pos}^{h}/t)+\sum_{i}\exp(s_{neg,i}^{h}/t)},$ (5)
where $t$ is a contrastive temperature hyperparameter.
@@
The proposed contrastive scoring metric quantifies the degree to which an attention head prioritizes the gold document for a given query relative to competing irrelevant documents. A higher value of $S_{CoRe}$ indicates that the corresponding attention head more effectively discriminates the positive document from hard negative documents.
Within the standard contrastive formulation, the temperature $t$ acts as a scaling factor that controls the sharpness of the attention score distribution between the positive and hard negative documents. A lower temperature produces a sharper distribution, imposing greater penalties on hard negatives and therefore identifying the most effective attention heads. However, excessively low temperatures may become overly selective, favoring only heads that perform exceptionally well at distinguishing the correct document while discarding heads that are broadly useful. Conversely, higher temperatures soften the distribution, reducing sensitivity to hard negatives and allowing for greater tolerance to documents that are semantically similar to the gold document.
@@
To identify CoRe heads, we compute the $S_{CoRe}$ of each attention head averaged over a set of data samples, and select the top-scoring few attention heads within each model as CoRe heads. By aggregating attention signals from these CoRe heads, the re-ranker achieves improved performance compared to aggregating over all heads. Figure 2 shows the affect of aggregating over different number of retrieval heads on the re-ranking results on the DBPedia dataset. We observe that using fewer – but higher quality – attention heads yields superior accuracy, reaching peak re-ranking accuracy with only $9$ CoRe heads.
@@
For each sample, we construct a LLM prompt consisting of the $50$ documents, followed by the instruction and query (see Appendix A for details). The positive document is inserted among the first five different positions, totaling to $5000$ samples per language model. We compute the $S_{CoRe}$ of each attention head on each sample, and compute the final retrieval head score by averaging across all samples. Less than top $1\%$ of all heads with the highest averaged score are selected as CoRe heads ($8$ heads in our experiment). This setup yields stable average retrieval scores, with the top CoRe heads consistently converging to the same subset for each model. Moreover, CoRe heads identified using a single dataset NQ generalize effectively across multiple datasets, enhancing re-ranking performance as shown later in Section 5.2.
@@
Once CoRe heads are identified for a given language model, re-ranking is performed by aggregating attention signals exclusively from this subset. Specifically, the new relevance score $s^{*}_{d_{i,j};q}$ in Equation 1 becomes:
$s^{*}_{d_{i,j};q}=\frac{1}{|\mathcal{T}_{q}|}\sum_{h\in H^{*}}\sum_{t\in\mathcal{T}_{q}}a_{j,t}^{h},$ (6)
where $H^{*}$ denotes the set of detected CoRe heads. The relevance score with respect to the content-free query $s^{*}_{d_{i,j};q_{cf}}$ is computed similarly. The calibrated relevance score $s^{*}_{d_{i,j}}$ of the $j$-th token in document $d_{i}$ is calculated as:
$s^{*}_{d_{i,j}}=s^{*}_{d_{i,j};q}-s^{*}_{d_{i,j};q_{cf}}.$ (7)
Finally, the calibrated relevance scores are aggregated to produce the document level re-ranking.
@@ 5 Experiments
In this section, we evaluate the effectiveness of CoRe-R across a diverse set of datasets and four different open-weight decoder models, comparing its performance against other zero-shot attention-based re-rankers. We also analyze the effect of layer pruning, showing CoRe-R achieves strong re-ranking performance without requiring the computation of all layers, significantly reducing computational overhead. Our code is available at https://github.com/linhhtran/CoRe-Reranking.
@@ 5.1 Setup
We compare CoRe-R with the following zero-shot attention-based re-rankers: ICR (Chen et al., 2025) and QR-R (Zhang et al., 2025). For QR head detection, we adopt the same detection data as described in Section 4 and exclude the contextual calibration step for a fair comparison with our head detection algorithm. Contextual calibration is applied only during the re-ranking stage, not the head detection step. For completeness, we report the re-ranking performance of QR-R with the original QR heads (detected with contextual calibration) in Section 5.4.
@@
We evaluate our approach on the BEIR benchmark (Thakur et al., 2021) which consists of fifteen diverse datasets: TREC-COVID (Voorhees et al., 2021), NFCorpus (Boteva et al., 2016), DBPedia-entity (Hasibi et al., 2017), SciFact (Wadden et al., 2020), SciDocs (Cohan et al., 2020), FiQA (Maia et al., 2018), NQ (Kwiatkowski et al., 2019), FEVER (Thorne et al., 2018), Climate-FEVER (Diggelmann et al., 2020), HotpotQA (Yang et al., 2018), Touche (Bondarenko et al., 2020), MSMARCO (Nguyen et al., 2016), Quora, ArguAna (Wachsmuth et al., 2018) and the CQADupStack series (Hoogeveen et al., 2015). Among them, NQ, HotpotQA, FEVER, Climate-FEVER, and DBPedia-entity are in-domain datasets and are derived from Wikipedia articles. Other datasets are out-of-domain. In addition, we evaluate on the Multilingual Long-Document Retrieval (MLDR) dataset (Chen et al., 2024) to assess cross-language generalization.
Re-ranking Details.
For each dataset in the BEIR benchmark, we re-rank the top-40 documents retrieved using the ibm-granite/granite-embedding-30m-english model (Awasthy et al., 2025). For the multilingual dataset MLDR, we re-rank the top-40 documents retrieved using the ibm-granite/granite-embedding-107m-multilingual model (Awasthy et al., 2025). Both QR-R and CoRe-R use the top $8$ retrieval heads, and report performance using the nDCG@10 scores. To ensure consistency, we adopt a uniform prompt structure for all datasets; details of the prompt design are provided in Appendix A.
@@
The temperature hyperparameter $t$ is tuned through grid search on a separate subset of the NQ train dataset. We select $t=0.001$ for Mistral 7B and Granite-3.2 8B, and $t=0.1$ for Llama-3.1 8B and Phi-4. These values are fixed for CoRe-R head detection prior to running the actual re-ranking experiments.

### v2

@@ 1 Introduction
To better characterize the functional role of attention heads, Wu et al. (2025) identify a subset of retrieval heads in transformers that specialize in extracting relevant information from long contexts. Their approach tracks copy–paste operations in Needle-in-a-Haystack tasks, which constrains its applicability to narrow question-answering scenarios. Building on this line of research, Zhang et al. (2025) propose QR heads (query-focused retrieval heads), selected according to the absolute attention allocated to the correct answer. While aggregated QR-head scores improve performance on downstream retrieval tasks, the selection criterion overlooks relative ranking. For instance, a head may assign high attention to the gold document yet allocate even greater weight to irrelevant documents, undermining its discriminative capacity. Since the QR-head method does not penalize such cases, it may fail to identify the heads that are most effective at distinguishing relevant from irrelevant information.
@@
  • We demonstrate that CoRe heads identified using a small subset of the Natural Questions (NQ) dataset generalize across the BEIR benchmark (Thakur et al., 2021) and even across multi-hop tasks, cross-lingual datasets and long context settings. Attention-based re-rankers using CoRe heads consistently outperform strong baselines, achieving state-of-the-art list-wise re-ranker performance.
  • We are the first to investigate layer pruning in attention-based re-rankers. We observe that top-scoring CoRe heads are primarily concentrated in the middle transformers layers. Exploiting this insight, we show that pruning most final layers incurs negligible loss in re-ranking accuracy, while reducing memory usage by $40\%$ and inference latency by $20\%$. This highlights the efficiency of CoRe heads for real-world retrieval systems.
@@ 2 Related Work
Zero-Shot Re-Ranking. Zero-shot re-ranking methods with LLMs can be grouped into three main approaches: point-wise, pair-wise, and list-wise. Point-wise methods (Sachan et al., 2022; Liang et al., 2022) score each document independently, typically through logits or direct generation. These approaches are computationally efficient but often underperform due to the absence of cross-document comparison. Pair-wise methods (Qin et al., 2024) compare two documents at a time and aggregate over all pairs to form a ranking. While improving accuracy, the computational cost grows quadratically with the number of documents. List-wise methods (Ma et al., 2023; Sun et al., 2023; Chen et al., 2025) consider all documents jointly to produce a single ranked list. Although list-wise methods require long-context modeling, advances in LLMs with extended context windows (Chen et al., 2023; Jin et al., 2024a; Fu et al., 2024) have made this increasingly practical. Consequently, list-wise approaches now achieve both scalability and strong effectiveness. Our work follows this trajectory, focusing on attention-based list-wise re-ranking (Chen et al., 2025), which offers more efficiency and stability compared to generation-based methods (Ma et al., 2023; Sun et al., 2023).
Role of Attention Heads. A growing body of work in mechanistic interpretability examines how the activation of attention heads shapes model behavior. Michel et al. (2019) and Voita et al. (2019) show that only a small fraction of heads are necessary for translation tasks. Olsson et al. (2022) identify induction heads that capture repeated input patterns, and later extended the work to heads supporting in-context learning (Yin and Steinhardt, 2025; Ren et al., 2024). Other studies reveal specialized heads responsible for knowledge conflict (Shi et al., 2024; Jin et al., 2024b) and context distraction (Zhu et al., 2025). More recently, Wu et al. (2025) identify retrieval heads, which copy answer tokens from long contexts into the output, while Zhang et al. (2025) propose query-focused retrieval heads (QR heads) that capture query–context interactions beyond direct copying. Collectively, these studies suggest that different subsets of heads specialize in distinct computational roles, and isolating the ones relevant to retrieval remains an open challenge. In this work, we build on the notion of retrieval heads (Wu et al., 2025; Zhang et al., 2025) as the mechanisms responsible for retrieving relevant information from long-context inputs – directly aligning with the re-ranking objective.
@@ 3 Background
Given a query $q$ and $k$ candidate documents $D=\{d_{1},d_{2},\dots,d_{k}\}$ returned by a base retriever, the goal of the re-ranker is to reorder $D$ so that the documents are sorted in the order of relevance to $q$ . To achieve this, Chen et al. (2025) propose In Context Re-ranking (ICR), an attention-based approach that leverages the attention signal from all attention heads of a LLM to compute fine-grained relevance scores.
Utilizing the long-context capability of LLMs, ICR constructs an input prompt consisting of a task instruction and a list of $k$ documents followed by the query $q$. Let $a_{j,t}^{h}$ be the attention score from the $t$-th token in query $q$ to the $j$-th token in document $d_{i}$ by an attention head $h$. For each token $j$ in document $d_{i}$, ICR computes the token-level relevance score with respect to the query $q$ within a head $h$ as
$s^{h}(d_{i,j})=\frac{1}{|\mathcal{T}_{q}|}\sum_{t\in\mathcal{T}_{q}}a_{j,t}^{h},$ (1)
where $\mathcal{T}_{q}$ denotes the set of the query tokens. We note that the defined token-level scores do not involve the attention scores of the instruction and separators, therefore they are not normalized post-softmax. The document-level relevance score of document $d_{i}$ within head $h$ is aggregated as:
$s^{h}(d_{i})=\sum_{t\in\mathcal{T}_{d_{i}}}s^{h}(d_{i,j})$ (2)
where $\mathcal{T}_{d_{i}}$ denotes the set of the document $d_{i}$’s tokens. As noted above, the token-level scores do not preserve the internal softmax of the attention head, hence the aggregated document-level scores are also not normalized. Aggregating over all heads in the head set $H$, the ICR document-level relevant scores is
$rel_{ICR}(d_{i})=\sum_{h\in H}s^{h}(d_{i}).$ (3)
@@
As noted above, ICR computes relevance scores by aggregating attention signals across all attention heads. While comprehensive, such full-head aggregation often introduces redundancy and noise, ultimately degrading re-ranking performance. Recent work on QR heads (Zhang et al., 2025) shows that leveraging only a targeted subset of heads $H_{QR}$ can improve retrieval effectiveness, suggesting that many attention heads are unnecessary for this task. Specifically, the QR document relevance score is computed as
$rel_{QR}(d_{i})=\sum_{h\in H_{QR}}s^{h}(d_{i}).$ (4)
The QR heads $H_{QR}$ is determined via a head detection process on a set of query-documents. Given a sample consisting of a query and $k$ candidate documents, QR rank a head $h$ based on its retrieval score
$S_{QR}(h)=s^{h}(d^{+})$ (5)
using the same notion of document-level score in Equation 2 for the positive document $d^{+}$.
The QR scoring metric essentially picks heads with the highest attention to the positive document. However, this metric does not account for the relative ranking within each head’s distribution as the document-level scores is not normalized. Because this criterion ignores how strongly a head contrasts the positive document against competing negatives, it fails to capture whether the head meaningfully separates signal from noise. This omission can lead to the selection of misleading heads. For example, as shown in Figure 2, selecting the top $8$ QR heads reduces accuracy on Quora for Mistral 7B and Llama 8B respectively, underperforming the ICR baseline that aggregates signals from all heads.
@@ 4 Contrastive Retrieval Heads
Addressing the limitations of prior attention-based re-ranking approaches, we propose a contrastive head detection framework that identifies a small subset of Contrastive Retrieval heads (CoRe heads) specialized for document re-ranking.
@@
Let $s^{h}(d_{i})$ denote the document-level score of document $d_{i}$ under attention head $h$ (Equation 2). We define the contrastive retrieval score of head $h$ as
$S_{CoRe}(h)=\frac{\exp(s^{h}(d^{+})/t)}{\exp(s^{h}(d^{+})/t)+\sum_{i\in D^{-}}\exp(s^{h}(d^{-}_{i})/t)},$ (6)
where $d^{+}$ is the positive document, $D^{-}$ denotes the set of negatives, and $t$ is a temperature hyperparameter. Intuitively, $S_{CoRe}(h)$ measures the extent to which head $h$ concentrates probability mass on the positive document relative to competing negatives. Heads with larger $S_{CoRe}$ therefore exhibit stronger discriminative ability and are more suitable for re-ranking.
Lemma 4.1 (Interpretation as head-level contrastive loss)
For a fixed query and candidate set $(d^{+},D^{-})$, ranking heads by $S_{CoRe}(h)$ is equivalent to ranking them by the negative log-likelihood
$\ell_{h}(s^{h})=-\log\Bigg(\frac{\exp(s^{h}(d^{+}))}{\exp(s^{h}(d^{+}))+\sum_{i\in D^{-}}\exp(s^{h}(d^{-}_{i}))}\Bigg).$ (7)
Lemma 4.1 shows that $S_{CoRe}$ admits a principled interpretation as a head-level contrastive objective, closely related to InfoNCE-style losses (Oord et al., 2018).
Comparison to QR-based Head Selection.
While the proposed contrastive scoring metric explicitly accounts for both positive and negative documents, the QR metric ranks heads solely based on the attention mass assigned to the positive document. Consequently, QR does not penalize heads that also assign substantial attention to irrelevant documents. The following proposition formalizes this limitation by showing that QR can select heads with strictly worse contrastive discrimination than an alternative favored by CoRe.
Proposition 4.2
(QR may select suboptimal heads under contrastive discrimination). Fix temperature $t>0$ and any candidate set containing one positive document $d^{+}$ and at least one negative document. There exist two attention heads $a$ and $b$ such that QR ranks $a$ above $b$ (i.e., $s^{a}(d^{+})>s^{b}(d^{+})$), yet $S_{CoRe}(a)<S_{CoRe}(b)$.
Proof. Consider the case with $k=2$ documents: one positive document $d^{+}$ and one negative document $d^{-}$. Let the document-level scores satisfy
$s^{a}(d^{+})>s^{b}(d^{+})\quad\text{and}\quad s^{a}(d^{-})\gg s^{a}(d^{+}),$
while
$s^{b}(d^{+})>s^{b}(d^{-}).$
Such configurations are feasible when document-level scores do not preserve the normalization post-softmax. In such cases, head $b$ is more preferable than head $a$.
Under QR, head $a$ is ranked above head $b$ since it assigns a larger score to the positive document. However, the contrastive score
$S_{CoRe}(h)=\frac{\exp(s^{h}(d^{+})/t)}{\exp(s^{h}(d^{+})/t)+\exp(s^{h}(d^{-})/t)}$
is decreasing in the margin $s^{h}(d^{-})-s^{h}(d^{+})$. Therefore, head $a$, which assigns more attention to the negative document than to the positive, satisfies $S_{CoRe}(a)<S_{CoRe}(b)$.
The same reasoning applies to any $k>2$, as $S_{CoRe}(h)$ depends monotonically on the total attention assigned to negative documents. Thus, QR may select heads with inferior positive–negative separation even in larger candidate sets.
@@
The proposed contrastive scoring metric quantifies the degree to which an attention head prioritizes the positive document for a given query relative to competing irrelevant documents. A higher value of $S_{CoRe}$ indicates that the corresponding attention head more effectively discriminates the positive from the negative documents.
Identifying CoRe heads.
@@
We compute the $S_{CoRe}$ of each attention head averaged over a small set of data samples, and select the top-scoring few attention heads within each model as CoRe heads. By aggregating attention signals from these CoRe heads, the re-ranker achieves improved performance compared to aggregating over all heads. Figure 2 shows the affect of aggregating over different number of retrieval heads on the re-ranking results on the DBPedia dataset. We observe that using fewer – but higher quality – attention heads yields superior accuracy, reaching peak re-ranking accuracy with only $9$ CoRe heads. Across models and datasets, we generally find that the optimal re-ranking performance is achieved with fewer than $10$ CoRe heads (see Appendix C.6 for details).
@@
For each sample, we construct a LLM prompt consisting of the $50$ documents, followed by the instruction and query (see Appendix A for details). The positive document is inserted among the first five different positions, totaling to $5000$ samples per language model. We compute the $S_{CoRe}$ of each attention head on each sample, and compute the final retrieval head score by averaging across all samples. Less than top $1\%$ of all heads with the highest averaged score are selected as CoRe heads ($8$ heads in our experiment). This setup yields stable average retrieval scores, with the top CoRe heads consistently converging to the same subset for each model. Moreover, CoRe heads identified using a single dataset NQ generalize effectively across multiple datasets, enhancing re-ranking performance as shown later in Section 5.2. We also note that the detection process is light-weight and relatively fast, requiring less than one hour on an H100 (96GB) GPU for the models considered in this work.
@@
Once the set of CoRe heads are identified for a given language model, re-ranking is performed by aggregating attention signals exclusively from this subset. Specifically, the CoRe document relevance score is
$rel_{CoRe}(d_{i})=\sum_{h\in H_{CoRe}}s^{h}(d_{i}),$ (8)
where $H_{CoRe}$ denotes the set of detected CoRe heads.
@@ 5 Experiments
In this section, we evaluate the effectiveness of CoRe heads across a diverse set of datasets and four different open-weight decoder models. We also analyze the effect of layer pruning, showing the proposed method achieves strong re-ranking performance without requiring the computation of all layers, significantly reducing computational overhead. Our code is available at https://github.com/linhhtran/CoRe-Reranking.
@@ 5.1 Setup
We compare CoRe-R with the existing zero-shot attention-based re-rankers: ICR (Chen et al., 2025) and QR-R (Zhang et al., 2025). For QR head detection, we adopt the same detection data as described in Section 4 for a fair comparison with our head detection algorithm. We also report comparison with the original retrieval heads (Wu et al., 2025) identified via copy-paste mechanism on Needle-in-a-Haystack task. We refer to these heads as NIAH heads and report the results in Section 5.4.
@@
We evaluate our approach on the BEIR benchmark (Thakur et al., 2021) which consists of fifteen diverse datasets: TREC-COVID, NFCorpus, DBPedia-entity, SciFact, SciDocs, FiQA, NQ, FEVER, Climate-FEVER, HotpotQA, Touche, MSMARCO, Quora, ArguAna and the CQADupStack series. Among them, NQ, HotpotQA, FEVER, Climate-FEVER, and DBPedia-entity are in-domain datasets, and other datasets are out-of-domain. In addition, we evaluate on the Multilingual Long-Document Retrieval (MLDR) dataset (Chen et al., 2024) for cross-language generalization and MuSiQue (Trivedi et al., 2022) for multi-hop generalization.
Re-ranking configuration.
For each dataset in the BEIR benchmark, we re-rank the top-$k$ documents retrieved using the ibm-granite/granite-embedding-30m-english model (Awasthy et al., 2025). For the multilingual dataset MLDR, we re-rank the top-$40$ documents retrieved using the ibm-granite/granite-embedding-107m-multilingual model (Awasthy et al., 2025). Following Chen et al. (2025), we re-rank MuSiQue top-$20$ retrieved from ColBERT v2. Both QR-R and CoRe-R use the top $8$ retrieval heads, and report performance using the nDCG@10 scores. To ensure consistency, we adopt a uniform prompt structure for all datasets; details of the prompt design are provided in Appendix A.
@@
The temperature hyperparameter $t$ is tuned through grid search on a separate subset of the NQ train dataset. We select $t=0.001$ for Mistral 7B and Granite-3.2 8B, and $t=0.1$ for Llama-3.1 8B and Phi-4. These values are fixed for CoRe-R head detection prior to running the actual re-ranking experiments. Note that the effect of temperature on CoRe head identification and subsequent re-ranking performance is examined in Section 5.5.